OpenStudy (anonymous):

(cos75+isin75)÷(cos15-isin15) simplify

OpenStudy (anonymous):

cos75 =sin15 ,sin 75=cos15

OpenStudy (anonymous):

simplify plz

OpenStudy (anonymous):

(sin 15+i cos 15)/(cos15-i sin15)

zepdrix (zepdrix):

Start by multiplying by the conjugate of the denominator, \[\Large\rm \frac{\cos75+\mathcal i \sin 75}{\cos15-\mathcal i \sin 15}\color{royalblue}{\left(\frac{\cos15+\mathcal i \sin 15}{\cos15+\mathcal i \sin 15}\right)}\] \[\Large\rm =\frac{(\cos75+\mathcal i \sin 75)(\cos15+\mathcal i \sin15)}{\cos^215+\sin^215}\]Apply our Pythagorean Identity for Sine and Cosine,\[\Large\rm =\frac{(\cos75+\mathcal i \sin 75)(\cos15+\mathcal i \sin15)}{1}\]So it turns into a multiplication problem now.\[\Large\rm =\cos(75+15)+\mathcal i \sin(75+15)\]I forget the name of the property that allows us to do that... if you convert to exponentials, the addition becomes a lot clearer.

OpenStudy (unklerhaukus):

Use Euler's formula: \[\Large\boxed{e^{i\theta}=\cos\theta+i\sin\theta}\] And some properties of (even) cosine and (odd) sine: \[\boxed{\cos\theta =\cos(-\theta)}\]\[\boxed{\sin\theta =-\sin(-\theta)}\] \[\frac{\cos(75°)+i\sin(75°)}{\cos(15°)-i\sin(15°)}\\ =\frac{\cos(75°)+i\sin(75°)}{\cos(-15°)+i\sin(-15°)}\\ =\frac{e^{75°i}}{e^{-15°i}}\\ =e^{(75°+15°)i}\\ %=e^{90°i}\\ %=\cos(90°)+i\sin(90°)\\ %=i =\dots \]